# Early Classification of Time Series: Taxonomy of Methods and Extensive Benchmark

### v2

@@ Abstract
In many situations, the measurements of a studied phenomenon are provided sequentially, and the prediction of its class needs to be made as early as possible so as not to incur too high a time penalty, but not too early and risk paying the cost of misclassification. This problem has been particularly studied in the case of time series, and is known as Early Classification of Time Series (ECTS). Although it has been the subject of a growing body of literature, there is still a lack of a systematic, shared evaluation protocol to compare the relative merits of the various existing methods. This document begins by situating these methods within a principle-based taxonomy. It defines dimensions for organizing their evaluation, and then reports the results of a very extensive set of experiments along these dimensions involving nine state-of-the art ECTS algorithms. In addition, these and other experiments can be carried out using an open-source library in which most of the existing ECTS algorithms have been implemented (see https://github.com/ML-EDM/ml_edm).
@@ 1 Introduction
In hospital emergency rooms (?), in the control rooms of national or international power grids (?), in government councils assessing critical situations, there is a time pressure to make early decisions. On the one hand, the longer a decision is delayed, the lower the risk of making the wrong decision, as knowledge of the problem increases with time. On the other hand, late decisions are generally more costly, if only because early decisions allow one to be better prepared. For example, a cyber-attack that is not detected quickly enough gives hackers time to exploit the security flaw found.
@@
A number of applications involve making decisions that optimizes a trade-off between accuracy of the prediction and its earliness. The problem is that favoring one usually works against the other. Greater accuracy comes at the price of waiting for more data.
Such a compromise between the Earliness and the Accuracy of decisions has been particularly studied in the field of Early Classification of Time Series (ECTS) (?), and introduced by ? (?). ECTS consists in finding the optimal time to trigger the class prediction of an input time series observed over time. Successive measurements provide more and more information about the incoming time series, and ECTS algorithms aim to optimize online the trade-off between two conflicting objectives, namely, the earliness and accuracy of class predictions. More formally, we have the following problem.
Problem statement:
In the ECTS problem, measurements of an input time series are observed over time. At time $t$, the incomplete time series ${\mathbf{x}}_{t}\,=\,\langle{x_{1}},\ldots,{x_{t}}\rangle$ is available where ${x_{i}}_{(1\leq i\leq t)}$ denotes the time indexed measurements. These measurements can be single or multi-valued. The input time series belongs to an unknown class $y\in{\cal Y}$. The task is to make a prediction $\hat{y}\in{\cal Y}$ about the class of the incoming time series, at a time $\hat{t}\in[1,T]$ before the deadline $T$ which corresponds to the length of the full time series. A misclassification cost is incurred when a prediction is made, denoted by $\mathrm{C}_{m}(\hat{y}|y)$. Furthermore, there exists a delay cost $\mathrm{C}_{d}(\hat{t})$ that expresses the time pressure and encourages early decisions (defined in Section 2.2). The choice of the best trigger time should optimize a compromise between these costs that move in opposite direction. The choice of the best triggering time must optimize a compromise between the two costs, which are moving in opposite directions.
To the best of our knowledge, ? (?) are the earliest explicitly mentioning “classification when only part of the series are presented to the classifier”. Since then, several researchers have continued their efforts in this direction and have published a large number of research articles. A recent and extensive review of the ETCS approaches can be found in the paper written by ? (?), including the applications that motivated the researchers to work in this area, and covering about fifty relevant papers selected from the $213$ papers found by search engines at the time of this writing.
As pointed out by ? (?), the ECTS problem is a special case of optimal stopping (?, ?), where the decision to be made concerns both: (i) when to stop receiving new measurements in order to (ii) predict the class of the incoming time series.
In the same paper, the ECTS problem has been extended into a more general one, which consists in optimizing the decision times of Machine Learning (ML) models in a wide range of settings where data is collected over time. The authors proposed a set of open research questions to the community, in order to widen the range of applications that are amenable to the ECTS framework (i.e. dealing with other learning tasks, other types of data, other application contexts etc.).
@@
However, despite the growing interest in ECTS over the last twenty years, there still remains a need for a shared taxonomy of approaches and an agreed well-grounded evaluation methodology. Here, in particular, we list limits that hamper a fair comparison of ECTS methods and algorithms:
@@
    Costs taken into account for evaluating the performance of the proposed method are not always clearly stated. It seems natural to distinguish between the misclassification costs $\mathrm{C}_{m}(\hat{y}|y)$, and the delay cost $\mathrm{C}_{d}(t)$, and to add them in order to define the cost of making a decision at time $t$. More generally, the delay cost may depend on the true class $y$ and the predicted one $\hat{y}$, and a single cost function $\mathrm{C}_{m}(\hat{y}|y,t)$ integrating misclassification and delay costs should then be used. For the sake of clarity, we keep the simple notation which distinguishes both cost functions in the rest of this paper. But in all cases, it is essential to state the framework used and the associated evaluation metric.
@@
    The performance of the proposed methods should be evaluated against a range of possible types of cost functions. It is usual to evaluate “by default” the methods using a $\ell_{0-1}$ loss function that penalizes a wrong classification by a unity cost, and to consider a linear delay cost function: $\mathrm{C}_{d}(t)=\lambda\,t$, for a value $\lambda>0$. However, lots of applications rather involve unbalanced miss-classification costs, and possibly also non linear delay costs. This is the case, for instance, in maintenance applications where wrongly not recognizing a critical situation is much more costly than wrongly predicting a problem and taking steps to fix it, and where delay cost may rise as a an exponential function of time: $\mathrm{C}_{d}(t)=\lambda\,e^{t}$ It is therefore quite important to assess the adaptability of the methods to various representative problem settings.
@@
    The contributions of the various components of a ECTS algorithm should be clearly delineated. The predominant approach to ECTS is to have a decision component which is in charge of evaluating the best moment to make the prediction about the class of the incoming times series, and a classifier one which makes the prediction itself. In order to fairly compare the triggering methods, which are at the heart of ECTS, the classifier used should be the same. We call these methods “separable methods”.
    An alternative approach relies on having a system that classifies the incoming time series $\mathbf{x}_{t}$ at each time with a prediction $\hat{y}$ in the set {‘postpone decision’, $y_{1},\ldots,y_{N}$} where $N$ is the number of classes. Therefore, within this approach, a single system decides either to wait at least one more time step, or to predict a class and stops the process. In this case, no distinction can be made between a decision component and a classifier one, and the whole system is evaluated as such. In the spirit of deep neural networks, we call this type of methods “end-to-end” to underline the fact that a single learning system is in charge of all operations, here decision and classification. Of course, this precludes a comparison involving only the choice of the decision component with other methods.
@@
    As with other supervised learning tasks, performance should be compared with that of “baseline” algorithms. In the case of ECTS tasks, two naive baselines are: (1) make a prediction as soon as it is allowed, and (2) make a prediction at $T$, after the entire time series has been observed. In our experiments reported in Section 4, we have added a third baseline, less simple than the two afore-mentioned ones, but still too obvious so that, to our knowledge, it has never been published as an original method. This is a confidence-based method where a decision is triggered as soon as the confidence for the likeliest class given $\mathbf{x}_{t}$ is greater than a threshold. (Formally, let $\hat{y}=\operatorname*{arg\,max}_{y\in\mathcal{Y}}p(y|\mathbf{x}_{t})$, then a prediction is made (i.e. $\hat{y}$) as soon as $p(\hat{y}|\mathbf{x}_{t})\geq\varepsilon$, for some threshold $\varepsilon\in[0,1]$.)
@@
    Precautions should be taken when using datasets of training time series to ensure that no bias enters unwillingly into the training and evaluation process. A case in point, concerns the normalization often used in time series datasets. ? (?) have reported that 71% of the reference time series classification datasets used to evaluate ECTS methods are made up of z-normalized time series, i.e. with measurements independently modified on each complete series to obtain a mean of zero and a standard deviation equal to 1. Clearly, this setting is not applicable in practice, as z-normalization would require knowledge of the entire incoming time series. In a research context, previous work has used such training sets to test the proposed algorithms. As ? (?, ?) note, this preprocessing is irreversible and can generate a problem for ECTS by introducing a temporal leakage of information. In order to assess its impact, we report in Section B.5 of Appendix B a comparison of results for z-normalized and non-normalized time series.
Up until now, it has been difficult to conduct fair comparisons between competing methods. Often, published performances are based on choices concerning data sets, the precise performance measure used, hyperparameter values and evaluation protocols (e.g., the split between training and test sets) that are not entirely explicit or, in any case, are difficult to reproduce. This is why we have recoded all the methods, specified a shared evaluation protocol with variants that can be employed by everyone, and searched for a collection of data sets that can be widely used to test and compare new as well as existing methods. We hope this will be a useful resource for the scientific community working in this field.
As this community shows a growing interest in the ECTS problem, granted by the increasing number of applications that fall in its range, it is timely (1) to propose a framework into which to cast the various approaches and thus indicate avenues for future research, and (2) a well-grounded evaluation methodology. Specifically, this paper makes the following contributions:
@@
  • A taxonomy is proposed in Section 2, classifying approaches in the literature according to their design choices.
@@
  • Extensive experiments have been performed, meeting the above-mentioned shortcomings. (1) The experimental protocol in Section 4.1 explicitly defines the costs used during training and evaluation, and varies the balance between misclassification and delay costs by using a large range of cost values. (2) Experiments are performed repeatedly for several types of cost function, i.e. balanced or unbalanced misclassification cost, and linear or exponential delay cost (see Sections 4.2 and 4.3) and many intermediate results are available in the supplementary materials. (3) Ablation and substitution studies are conducted in Section 4.4 with the aim of evaluating the impact of methodological choices, such as the choice of classifier, its calibration, or even z-normalization of training time series. (4) The experiments include three baseline approaches, rarely considered in the literature, which often give surprisingly good results. (5) In addition to the reference data used in the ECTS field, a collection of some thirty non-z-normalized datasets is proposed and provided to the community.
@@
  • An open source library is being made available [^1] https://github.com/ML-EDM/ml_edm in order to enable reproducible experiments, as well as facilitate the scientific community’s development of future approaches. Particular care has been taken to ensure the quality of the code, so that this library may be used to develop real-life applications.
The rest of this paper is organised as follows : The section 2 proposes and describes a new ECTS taxonomy, the different choices to be made in a well-founded way when designing an ECTS method and a set of four questions that need to be answered to make these choices.
Section 3 presents a comprehensive view of the ECTS field, along the suggested taxonomy and the four questions raised in Section 2. In Section 4 we present the pipeline developed in order to realize extensive experimentation and we report the main results obtained for different cost setting. This benchmark is supported by a library released as Open Source for dissemination and used in the ECTS research community. Finally Section 5 concludes this paper. Appendix A lists the datasets used for the experiments, and complementary results are provided in Appendix B.
2 Organizing the ECTS approaches: a taxonomy
@@
The aim of this section is to outline in a principled way the various choices that need to be made when designing one ECTS method.
General form of an ECTS model
@@
  • $\mathrm{C}_{m}(\hat{y}|y):{\cal Y}\times{\cal Y}\rightarrow\mathbb{R}$, that corresponds to the misclassification cost of predicting $\hat{y}$ when the true class is $y$.
@@
$t^{\star}=\operatorname*{arg\,min}_{t\in[1,T]}\mathcal{L}(\hat{y}({\mathbf{x}}_{t}),t,y).$ (1)
@@
$s^{\star}({\mathbf{x}}_{t})=\left\{\begin{array}[]{ll}\emptyset&\mbox{if extrameasures are queried;}\\ y^{\star}=\hat{y}(x_{t^{\star}})&\mbox{when prediction is triggered at $t=t^{\star}$;}\end{array}\right.$ (2)
ECTS is however an online optimization problem, where at each time step $t$ a function $s({\mathbf{x}}_{t})$ must decide to make a prediction or not. Equation 1 is thus no longer operational since it requires a complete knowledge of the time series. In practice, the function $s({\mathbf{x}}_{t})$ triggers a decision at $\hat{t}$, based on a partial description ${\mathbf{x}}_{\hat{t}}$ of the incoming time series ${\mathbf{x}}_{T}$ (with $t\leq T$). The goal of an ECTS system is to chose a triggering time $\hat{t}$ as close as possible to the optimal one $t^{*}$, at least in terms of cost, minimizing $\mathcal{L}(\hat{y}({\mathbf{x}}_{\hat{t}}),\hat{t},y)-\mathcal{L}(\hat{y}({\mathbf{x}}_{t^{\star}}),t^{\star},y)$ as much as possible.
From a machine learning point of view, the goal is to find a function $s\in{\cal S}$ that best optimizes the loss function $\mathcal{L}$, minimizing the true risk over all time series distributed according to the distribution222 Notice that the notation $\mathcal{X}$ is an abuse that we use use to simplify our purpose. In all mathematical rigor, the measurements observed successively constitute a family of time-indexed random variables $\mathbf{x}=(\mathbf{x}_{t})_{t\in[1,T]}$. This stochastic process $\mathbf{x}$ is not generated as commonly by a distribution, but by a filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\in[1,T]}$ which is defined as a collection of nested $\sigma$-algebras (?) allowing to consider time dependencies. Therefore, the distribution $\mathcal{X}$ should also be re-written as a filtration. $\mathbb{P}_{\mathcal{X}}$ that governs the time series in the application:
$\small\normalsize\operatorname*{arg\,min}_{s\in\mathcal{S}}\operatornamewithlimits{\mathbb{E}_{\>\mathbf{x}\sim\mathbb{P}_{\mathcal{X}}}^{\>}}\left[\mathcal{L}(\hat{y}({\mathbf{x}}_{\hat{t}}),\hat{t},y)\right]$ (3)
The questions then are:
  1. 1.
    Which form can take the function $s(\cdot)$? We will distinguish end-to-end architecture from separable ones.
  2. 2.
    How the criterion to be optimized accounts for the trade-off between accuracy and earliness? We will see that the costs implied, about misclassification and delay, are variously explicit in the existing methods.
  3. 3.
    How the when question of the stopping problem can be approached? This will lead us to distinguish between cost-informed and cost-uninformed ones on one hand, and between anticipation-based methods versus myopic ones, on the other hand.
  4. 4.
    How the prediction problem itself can be solved given that $\mathbf{x}_{t}$ belongs to a different input spaces at each time step $t$?
This set of questions and possible choices for their solution are illustrated in Figure 1. We turn successively to each one in the following.
Figure 1: Proposed ECTS taxonomy
2.1 The different forms of the function $s(\cdot)$
@@
In the separable approach, these questions are solved using two separate components. The classification one deals with making a prediction: $\mathbf{x}_{t}\mapsto\hat{y}$, while the trigger function decides when to predict. Within this perspective, the classification component is learned independently of the trigger one, while the latter uses the results of the classification component in order to trigger a decision. A simple triggering strategy is to decide it is time to make a prediction as soon as the classification component is sufficiently sure of its prediction. We formalize separable approaches by: $s(\mathbf{x}_{t})=(g\circ h)(\mathbf{x}_{t})$ where $g$ is the decision or trigger function, and $h$ is the prediction function.
@@
In the end-to-end approaches, a single component decides when to make a prediction and what that prediction is. Thus, the function $s$, defined in Equation 2, is responsible both for choosing the time $\hat{t}$ for making the prediction, and for the prediction itself $\hat{y}$.
@@
The question that naturally arises is which type of architecture (i.e. end-to-end or separable) performs best. On the one hand, in separable approaches, the classification component is trained independently of the triggering one, which can be detrimental. On the other hand, separating the ECTS problem into two inherently simpler sub-problems could be an advantage. In this paper, we do not delve any further into this question, which we leave for future work.
2.2 Choice of the optimizing criterion during training
This section covers optimization criteria existing in the literature and used to train ECTS functions. In the following, these criteria are listed by level of cost awareness, and we differentiate between the following situations:
  1. 1.
    The first one, we call cost-informed at training time.
    $\mathbb{P}_{\mathcal{X}}$ being unknown, instead of using Equation 3 describing the true risk, one tries to minimize the empirical risk, also called average cost in the ECTS literature, for a training set of $M$ time series:
    $AvgCost\;=\frac{1}{M}\sum_{i=1}^{M}\mathcal{L}(\hat{y}_{i},y_{i},\hat{t})\;=\;\frac{1}{M}\sum_{i=1}^{M}\mathrm{C}_{m}(\hat{y}_{i}|y_{i})+\mathrm{C}_{d}(\hat{t}_{i})$ (4)
    AvgCost is the most appropriate criterion to both train and evaluate ECTS approches. The following presents proxy measures of AvgCost.
  2. 2.
    The second situation is a sub-case of cost-informed, and can be qualified as cost-proxy at training time. There, while the accuracy and earliness of prediction are taken into account, the optimization criterion combines them in a proxy which is not the AvgCost. Classical proxys include:
    • The Harmonic Mean (see ?):
      $HM=\frac{2\times Accuracy\times(1-Earliness)}{Accuracy+(1-Earliness)}$ (5)
      with:
      $Accuracy$ $=\frac{1}{M}\sum_{i=1}^{M}\mathbb{1}(\hat{y}_{i}=y_{i})$ (6)
      $Earliness$ $=\frac{1}{M\times T}\sum_{i=1}^{M}\hat{t}_{i}$ (7)
      where $\hat{t}_{i}$ is the time that the ECTS function decides to make prediction for the time series $\mathbf{x}^{i}$.
    • The CF (i.e. Cost Function) criterion with $0\leq\alpha\leq 1$ (see ?, ?, ?):
      $CF=\alpha\times(1-Accuracy)+(1-\alpha)\times Earliness$ (8)
      When the cost of misclassification $\mathrm{C}_{m}(\hat{y},y)=\mathbb{1}(\hat{y}=y)$, and the delay cost is $\mathrm{C}_{d}(t)=\hat{t}/T$ and $\alpha=0.5$, CF becomes a particular case of AvgCost.
    The choice of using costs or approximating them by a proxy is a technical issue, which may, for example, be relevant to making the loss function differentiable by approximation. For this reason, in the remainder of this paper, cost-informed/cost-proxys at training time situations are grouped together under the term cost-informed at training time, this difference not being essential.
  3. 3.
    The third situation is qualified as cost-uninformed-train. This is the case of methods that only set the threshold on the confidence of the prediction in order to trigger the prediction, regardless of the costs incurred on the training set. Another example is methods that use rigid rules, such as “decide as early as the first measurement is available”.
It is a question whether any of these approaches fare better than the others. To measure this, we must use the Average Cost (see Equation 4) measured on a test set, which represents the ground truth of the ECTS problem. It should be noted that many of the proposed methods have been evaluated on the basis of other criteria in the literature, with the result of not allowing a rigorous comparison between them. We will come to this problem in Section 4.
@@ 2.2 Choice of the optimizing criterion during training
The rest of this section is specific to separable ECTS approaches, which represent a large part of the literature.
2.3 Information used by the trigger function during inference
The trigger function can draw on different types of information. In the simplest case, it can decide irrespectively of the incoming time series $\mathbf{x}_{t}$. This is the case, for example, of the rule that would say: “wait until half of the measurements are available, then make a prediction”. The corresponding trigger function can be said to be blind.
Apart from this extreme case, it is interesting to distinguish two dimensions. First, the trigger function may or may not take the misclassification and delay costs into account. Second, it can also make its decision at each time step solely on the basis of past observations, or it can anticipate possible futures to help its decision.
For instance, confidence-based methods (seen Section 2.3.1) do not explicitly take costs into account in the triggering decision.

### v6

# Early Classification of Time Series: A Survey and Benchmark

@@ Abstract
In many situations, the measurements of a studied phenomenon are provided sequentially, and the prediction of its class needs to be made as early as possible so as not to incur too high a time penalty, but not too early and risk paying the cost of misclassification. This problem has been particularly studied in the case of time series, and is known as Early Classification of Time Series (ECTS). Although it has been the subject of a growing body of literature, there is still a lack of a systematic, shared evaluation protocol to compare the relative merits of the various existing methods. In this paper, we highlight the two components of an ECTS system: decision and prediction, and focus on the approaches that separate them. This document begins by situating these methods within a principle-based taxonomy. It defines dimensions for organizing their evaluation and then reports the results of a very extensive set of experiments along these dimensions involving nine state-of-the-art ECTS algorithms. In addition, these and other experiments can be carried out using an open-source library in which most of the existing ECTS algorithms have been implemented (see https://github.com/ML-EDM/ml_edm).
@@ 1 Introduction
In hospital emergency rooms (Mathukia et al., 2015), in the control rooms of national or international power grids (Dachraoui et al., 2015), in government councils assessing critical situations, in many situations there is a time pressure to make early decisions. On the one hand, the longer a decision is delayed, the lower the risk of making the wrong decision, as knowledge of the problem increases with time. On the other hand, late decisions are generally more costly, if only because early decisions allow one to be better prepared.
@@
A number of applications thus involve making decisions that optimize a trade-off between the accuracy of the prediction and its earliness. The problem is that favoring one usually works against the other. Greater accuracy comes at the price of waiting for more data. Such a compromise between the Earliness and the Accuracy of decisions has been particularly studied in the field of Early Classification of Time Series (ECTS), and introduced by Xing et al. (2008). ECTS consists in finding the optimal time to trigger the class prediction of an input time series observed over time.
As pointed out by Bondu et al. (2022), the ECTS problem is rooted in optimal stopping, of which it is a special case (Shepp, 1969; Ferguson, 1989), where the decision to be made concerns both: (i) when to stop receiving new measurements in order to (ii) predict the class of the incoming time series. To the best of our knowledge, Alonso González & Diez (2004) are the earliest explicitly mentioning “classification when only part of the series are presented to the classifier”. Since then, several researchers have continued their efforts in this direction and have published a large number of research articles (Xing et al., 2009; Anderson et al., 2012; Dachraoui et al., 2015; Mori et al., 2017a; Schäfer & Leser, 2020; Lv & Hu, 2022; Ebihara et al., 2025).
@@
However, despite the growing interest in ECTS over the last twenty years (Gupta et al., 2020; Akasiadis et al., 2024; Santos & Kern, 2016), there still remains a need for a shared taxonomy of approaches and an agreed well-grounded evaluation methodology (see Table 3 in Appendix A). Guidelines that we feel are important for a fair and informative comparison of existing ECTS approaches are listed below:
@@
    Costs taken into account for evaluating the performance of the proposed method should be explicitly stated. It seems natural to distinguish between the misclassification costs, and the delay cost, and to add them in order to define the cost of making a decision at time $t$. The delay cost may also depend on the true class $y$ and the predicted one $\hat{y}$, and a single cost function integrating misclassification and delay costs should then be used. For the sake of clarity, we keep the simple notation that distinguishes both cost functions in the rest of this paper.
@@
    The performance of the proposed methods should be evaluated against a range of possible types of cost functions. It is usual to evaluate “by default” the methods using a $\ell_{0-1}$ loss function that penalizes a wrong classification by a unity cost, and to consider a linear delay cost function. However, lots of applications rather involve unbalanced misclassification costs, and possibly also non-linear delay costs.
@@
    The contributions of the various components of an ECTS algorithm should be as clearly delineated as possible. The predominant approach to ECTS is to have a decision component which is in charge of evaluating the best moment to make the prediction about the class of the incoming time series, and a classifier one which makes the prediction itself. We call these methods “separable methods”. In order to fairly compare the triggering methods at play, which are at the heart of ECTS, the classifier used should be the same for all methods in the experiments.
    Recent approaches mix the decision and the prediction components. This is the case of “end-to-end” methods using neural networks. These methods do not allow the evaluation of the merits of the decision component by itself, independently of the classification component, and they are accordingly not considered here.
@@
    Performance obtained should be compared with that of “baseline” algorithms. Failing to do this weakens any claim about the value of the proposed method. In the case of ECTS tasks, two naive baselines are: (1) make a prediction as soon as it is allowed, and (2) make a prediction after the entire time series has been observed. In our experiments reported in Section 4, we have added a third baseline less simple than the two aforementioned, which, to our knowledge has never been published as an original method. This is a confidence-based method where a decision is triggered as soon as the confidence for the likeliest class given $\mathbf{x}_{t}$ is greater than a threshold.
@@
    The datasets used for training and testing should remove the biases that are specific to the ECTS task and that may result in erroneous evaluations. A case in point, concerns the normalization often used in time series datasets. Dau et al. (2019) have reported that 71% of the reference time series classification datasets used to evaluate ECTS methods are made up of z-normalized time series, Not only, this setting is not applicable in practice, as z-normalization would require knowledge of the entire incoming time series, but it could also introduce a leakage of information from the future of the time series to the previous time steps which is detrimental to a fair assessment of the methods (Wu et al., 2021).
In this article, we aim to present a methodology for a fair and informative comparison between separable ECTS methods in the literature. Specifically, this paper makes the following contributions:
@@
  • A taxonomy is proposed in Section 2, classifying separable approaches in the literature according to their design choices.
@@
  • Extensive experiments have been performed, taking into account the obstacles mentioned above. (1) The experimental protocol in Section 4.1 explicitly defines the costs used during training and evaluation, and varies the balance between misclassification and delay costs by using a large range of cost values. (2) Experiments are performed repeatedly for several types of cost function, i.e. balanced or unbalanced misclassification costs, and linear or exponential delay costs (see Sections 4.2 and 4.3) and many intermediate results are available in the supplementary materials. (3) Ablation and substitution studies are conducted in Section 4.4 with the aim of evaluating the impact of methodological choices, such as the choice of classifier, its calibration, or even z-normalization of training time series, as well as the non-myopia property of some trigger functions. (4) The experiments include three baseline approaches, rarely considered in the literature, which often prove to be efficient. (5) In addition to the reference data used in the ECTS field, a collection of some thirty non-z-normalized datasets is proposed and provided to the community.
@@
  • An open source library is made available111(see https://github.com/ML-EDM/ml_edm) to enable reproducible experiments, as well as to facilitate the scientific community’s development of future approaches. Particular care has been taken to ensure the quality of the code, so that this library may be used to develop real-life applications (see Appendix D).
The scope of this paper is a survey and a benchmark of separable ECTS approaches, in line with our guideline #3. Interested readers can refer to Appendix C for complementary results on end-to-end recent approaches. The rest of this paper is organized as follows: Section 2 proposes and describes an ECTS taxonomy in order to situate separable approaches in relation to each other. It underlines the different choices to be made in a well-founded way when designing an ECTS method. Section 3 presents a review of state-of-the-art methods for ECTS organized along the dimensions of the suggested taxonomy. In Section 4, we present the pipeline developed to conduct extensive experimentation, and we report the main results obtained for different cost settings. Finally, Section 5 concludes this paper.
Position with respect to other literature surveys
In reaction to the abundant scientific activity around ECTS, Gupta et al. (2020) wrote a survey, however with the following characteristics. First, their perspective on ECTS is confidence-based as they declare “A primary task of an early classification approach is to classify an incomplete time series as soon as possible with some desired level of accuracy” (italics are ours) and it is not centered on optimizing the tradeoff between predictive accuracy and earliness. Second, they organize the covered approaches by application domains and methods for representing time series, viz. prefix-based, shapelet-based, model-based, and miscellaneous. A perspective that, in a way, is tangential to the problem of selecting the right decision time.
Another paper (Akasiadis et al., 2024) intends to propose a framework to evaluate ECTS systems. Noticeably, it also adopts a confidence-based approach: “The objective is to find the earliest time-point of a time-series at which a reliable prediction can be made”. Alongside the survey, a benchmark is proposed whose characteristics are detailed in Table 3.
The earliest survey published in 2016 by Santos & Kern (2016) deals mostly with classification methods for complete time series. The final section describes four ECTS algorithms with no experiments.
2 ECTS: concepts and taxonomy
@@
The aim of this section is (i) to formally describe the ECTS problem, and (ii) to outline in a principled way the various choices that need to be made when designing one ECTS method.
Problem statement
In the ECTS problem, an input time series of size $T$ is progressively observed over time. At time $t\leq T$, the incomplete time series ${\mathbf{x}}_{t}\,=\,\langle{x_{1}},\ldots,{x_{t}}\rangle$ is available where ${x_{i}}_{(1\leq i\leq t)}$ denotes the time-indexed measurements. These measurements can be single or multi-valued. The input time series belongs to an unknown class $y\in{\cal Y}$. The task is to make a prediction $\hat{y}\in{\cal Y}$ about the class of the incoming time series, at a time $\hat{t}\in[1,T]$ before the deadline $T$.
@@
  • $\mathrm{C}_{m}(\hat{y}|y):{\cal Y}\times{\cal Y}\rightarrow\mathbb{R}$, that corresponds to the cost of predicting $\hat{y}$ when the true class is $y$.
@@
$t^{\star}=\operatorname*{arg\,min}_{t\in[1,T]}\mathcal{L}(\hat{y}({\mathbf{x}}_{t}),y,t).$ (1)
@@
[…]
ECTS is however an online optimization problem, where at each time step $t$ a function $s({\mathbf{x}}_{t})$ must decide whether to make a prediction or not. Equation 1 is thus no longer operational since it requires complete knowledge of the time series. In practice, the function $s({\mathbf{x}}_{t})$ triggers a decision at $\hat{t}$, based on a partial description $^{\mathbf{x}}_{\hat{t}}$ of the incoming time series ${\mathbf{x}}_{T}$ (with $\hat{t}\leq T$). The goal of an ECTS system is to choose a triggering time $\hat{t}$ as close as possible to the optimal one $t^{*}$, at least in terms of cost, minimizing $\mathcal{L}(\hat{y}({\mathbf{x}}_{\hat{t}}),y,\hat{t})-\mathcal{L}(\hat{y}({\mathbf{x}}_{t^{\star}}),y,t^{\star})$ as much as possible.
From a machine learning point of view, the goal is to find a function $s\in{\cal S}$ that best optimizes the loss function $\mathcal{L}$, minimizing the true risk over all time series distributed according to the distribution [^2] Notice that the notation $\mathcal{X}$ is an abuse that we use use to simplify our purpose. In all mathematical rigor, the measurements observed successively constitute a family of time-indexed random variables $\mathbf{x}=(\mathbf{x}_{t})_{t\in[1,T]}$. This stochastic process $\mathbf{x}$ is not generated as commonly by a distribution, but by a filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\in[1,T]}$ which is defined as a collection of nested $\sigma$-algebras (Klenke, 2013) allowing to consider time dependencies. Therefore, the distribution $\mathcal{X}$ should also be re-written as a filtration. $\mathbb{P}_{(\mathcal{X}\times\mathcal{Y})}$ that governs the time series in the application:
$\operatorname*{arg\,min}_{s\in\mathcal{S}}\mathbb{E}_{(\mathbf{x},y)\sim\mathbb{P}_{(\mathcal{X}\times\mathcal{Y})}}\left[{\mathcal{L}(\hat{y}({\mathbf{x}}_{\hat{t}}),y,\hat{t})}\right$] (3)
In order to solve the ECTS problem, a training set composed of $M$ labeled time series, denoted by $({\mathbf{x}}^{i}_{T},y^{i})_{i\in[0,M]}\in(\mathcal{X}\times\mathcal{Y})$, where each series ${\mathbf{x}}_{T}=\,\langle{x_{1}},\ldots,{x_{T}}\rangle$ is complete and of the same size $T$, and associated with its label $y\in\mathcal{Y}$, is used to learn how to predict the class of an incoming time series ${\mathbf{x}}_{t}\,=\,\langle{x_{1}},\ldots,{x_{t}}\rangle$, and to learn what can be expected in the future given ${\mathbf{x}}_{t}$.
Consequently, model training and deployment are of different natures. The training stage is carried out as a supervised batch process, with access to the full labeled time series. When it comes to testing, on the other hand, decision-triggering is an online process which stops at time $\hat{t}$, and at the latest, when the deadline $T$ is reached.
2.1 Separable vs. end-to-end approaches
@@
In the separable approach, these questions are solved using two separate components. The classification one deals with making a prediction: $\mathbf{x}_{t}\mapsto\hat{y}$, while the trigger function decides when to predict. Within this perspective, the classification component is learned independently of the trigger one, while the latter uses the results of the classification component in order to trigger a decision. We formalize separable approaches by:
$s(\mathbf{x}_{t})=(g\circ h)(\mathbf{x}_{t})$ (4)
where $g$ is the decision or trigger function, and $h$ is the prediction function.
@@
In the end-to-end approaches, a single component decides when to make a prediction and what that prediction is. Thus, the function $s$, defined in Equation 2, is responsible both for choosing the time $\hat{t}$ for making the predictions, and for the prediction itself $\hat{y}$.
@@
The question that naturally arises is which type of architecture (i.e. end-to-end or separable) performs best. On the one hand, in separable approaches, the classification component is trained independently of the triggering one, which can be detrimental, for example, by propagating errors from one module to another. On the other hand, separating the ECTS problem into two inherently simpler sub-problems could be an advantage, for example, in terms of convergence during training. Additionally, the separable framework allows one to directly leverage the latest advances in the Time Series Classification (TSC) literature (Bagnall et al., 2017). In this paper, we do not delve any further into the question of which type of architecture is best, which we leave for future work.
@@
The rest of this section is specific to separable ECTS approaches, which represent a large part of the literature. In the following, we first examine the decision component (i.e. the trigger function) and then the prediction one (i.e. the classifier). Figure 1 gives a general view of the proposed taxonomy.
2.2 Trigger function’s properties
[…]
Figure 1: Proposed taxonomy for separable ECTS approaches, which can be viewed as a function composition of trigger and classification functions. The taxonomy does not follow a tree structure as the trigger function may have several properties, that are not mutually exclusive. For instance, a trigger function can be both confidence-based and anticipation based.
2.2.1 Cost-informed or cost-uniformed
The first property is whether or not the trigger function actually takes as input any cost functions (or eventually some proxy of them) during training. We call these approaches cost-informed. On the contrary, cost-uninformed approaches are usually based on some hard pre-defined rules to trigger and are not always easily expandable to a generic cost-setting framework. Additionally, when being cost-informed, trigger functions can leverage cost information even more, by being adaptable at testing time, i.e. during inference, trigger decisions directly depend on cost functions.